OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 3(2x - 1) greater than or equal to 4(2x - 3) - 3? A. x less than or equal to B. x greater than or equal to C. A number line with closed circle on 6 shading to the left. D. A number line with closed on 6 shading to the right.

OpenStudy (anonymous):

To solve the inequality, simplify the expressions first 3(2x - 1) => 4(2x - 3) - 3

OpenStudy (campbell_st):

1st thing to do is solve the inequality and see what you have as a solution

OpenStudy (anonymous):

Distribute 3 to 2x - 1. On the right side of the inequality, distribute 4 to 2x - 3.

OpenStudy (anonymous):

Can I do distributed property first?

OpenStudy (anonymous):

6x-1=>8x-12-12?

OpenStudy (anonymous):

Do like terms now ?

OpenStudy (anonymous):

Simplify the right side of the inequality further!

OpenStudy (anonymous):

6x-1=>8x?

OpenStudy (anonymous):

Also, on the right side, it's -12 - 3.

OpenStudy (anonymous):

I canceled 12 out

OpenStudy (anonymous):

So that turns into 8x - 15.

OpenStudy (anonymous):

So 6x-1=>8x-15? Now combine like terms right?

OpenStudy (anonymous):

It's also 6x-3. -1 * 3 = 3. The equation now looks like 6x - 3 => 8x - 15 Now, solve the inequality by isolating x. Hint! Subtract 6x from both sides of the equation.

OpenStudy (anonymous):

2x-3=>-15 add 3 to both sides right? and I would get -12?

OpenStudy (anonymous):

Okay so the equation is −3≥8x−15−6x. It simplifies into −3≥2x−15. Now, add 15 to both sides to get 2x by itself.

OpenStudy (anonymous):

\[2x \ge 12 \]

OpenStudy (anonymous):

Thanks you I got \[x \le 6\]

OpenStudy (anonymous):

Yep that's correct